OpenStudy (anonymous):

simplify the radical expression. 7. √63x^15y^9/7xy^11 8. √3+4√3 5√6 5√3 3√6 3√3 9. 2√6+3√96 14√6 14√96 5√96 50√6 10. (8 +√11 )(8 –√11 ) 53 75 + 16√11 –57 64 + √11

OpenStudy (pfenn1):

\[\sqrt{\frac{63x^{15}y^9}{7xy^{11}}}\]Let's try to simplify the expression.\[=\sqrt{\left( 63 \over7 \right)\left( x^{15}\over x \right)\left( y^9 \over y^{11} \right)}=\] can you do it from here?

OpenStudy (anonymous):

9X^7/Y

OpenStudy (anonymous):

is it right

OpenStudy (pfenn1):

you forgot to take the square root of 9

OpenStudy (anonymous):

3x^7/y

OpenStudy (pfenn1):

Great!

OpenStudy (anonymous):

thanks

OpenStudy (pfenn1):

you're welcome.